# Thermodynamic computing out of equilibrium

### v2

@@ II.1 Equilibrium behavior
$\langle\cdot\rangle_{0}\equiv\frac{\int{\rm d}x\,(\cdot)\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}{\int{\rm d}x\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}.$ (3)
When the neuron potential is purely quadratic, i.e. ${\bm{J}}=(J_{2},0,0)$ – with $J_{2}>0$ to ensure thermodynamic stability – the integrals in (3) can be solved analytically, giving the linear form $m=I/(2J_{2})$. This form is plotted, for $J_{2}=1$, as a gray line in Fig. 1(b); the horizontal dotted black line denotes the value zero. In this case the equilibrium activation function of the neuron is linear, meaning that networks of such neurons in equilibrium cannot serve as universal function approximators  [^2] The quadratic nonlinearity also gives rise to a linear activation function out of equilibrium..
Figure 2: (a) Equilibrium fluctuations (4) of the thermodynamic neuron of Fig. 1(a), for the case $\beta=100$. The addition of the quadratic coupling to the quartic one (green), suppresses fluctuations relative to the pure quartic case (blue). The mean equilibrium activation functions of those two cases are similar; see Fig. 1(c). The inset shows the largest fluctuations of the purely quartic neuron to be many times that of the purely quadratic neuron. (b,c) The mean (b) and variance (c) of the equilibrium activation function of the $(1,0,1)$ neuron depend on temperature.
Figure 3: Nonequilibrium properties of the thermodynamic neuron of Fig. 1(a). (a–c) Langevin evolution (5) of the neuron $x$ for $\beta=100$, for 11 evenly-spaced values of $I$. We consider three different sets of couplings ${\bm{J}}$: (a) purely quadratic; (b) purely quartic; and (c) mixed quadratic-quartic. (d) The finite-time activation function of the (1,0,1) neuron under the dynamics (5) (here for $\beta=\infty$) is nonlinear above some threshold observation time (here about 0.2 time units), and for longer times converges to the equilibrium result (curved black dashed line).
The simplest case that is thermodynamically stable and admits a nonlinear equilibrium activation function is the purely quartic case, ${\bm{J}}=(0,0,J_{4})$, with $J_{4}>0$. This case is shown as a blue line in Fig. 1(b), with $J_{4}=1$  [^3] In the purely quartic case the equilibrium activation function can be expressed analytically, in terms of the hypergeometric function.. The equilibrium activation function is nonlinear: its gradient (shown inset) is largest near the origin, and decreases as $|I|$ becomes large. The gradient remains finite for finite $I$: the neuron does not saturate.
Panels (c), (d), and (e) of Fig. 1 show the effect on the thermodynamic neuron’s equilibrium activation function of including quadratic and cubic terms with the quartic coupling. The resulting activation functions display a range of forms that resemble some of those used in classical neural networks, notably the sigmoid and hyperbolic tangent functions LeCun et al. (2015); Schmidhuber (2015); Dubey et al. (2022).
@@
$\sigma^{2}=\langle x^{2}\rangle_{0}-\langle x\rangle_{0}^{2}.$ (4)
In Fig. 2(a) we plot the value of (4) as a function of neuron input $I$ for the quadratic case ${\bm{J}}=(1,0,0)$ (gray) and the quartic case ${\bm{J}}=(0,0,1)$ (blue). In the quadratic case the variance is constant for constant temperature, $\sigma^{2}=1/(2\beta)$, reflecting the equipartition theorem. In the quartic case the variance is not constant, and is largest near $I=0$; the inset shows that its maximum value is several times that of the quadratic neuron. The mixed quadratic-quartic case ${\bm{J}}=(1,0,1)$, shown green in Fig. 1(c), has fluctuations at the origin comparable to the quadratic case, while the mean activation function of the case $(1,0,1)$, shown in Fig. 1(c), is similar to that of the pure quartic case. The addition of the quadratic term to the quartic one suppresses fluctuations without changing the essence of the nonlinearity.
@@
The equilibrium activation function of a nonlinear thermodynamic neuron depends on temperature. In Fig. 2(b) and (c) we show the mean and variance of the equilibrium activation function of the neuron ${\bm{J}}=(1,0,1)$ for a range of values of $\beta$. The mean activation is nonlinear provided that temperature is not much larger than the scales of $J_{2}$ and $J_{4}$. The case in which they are comparable, i.e. $|J_{2}|/k_{\rm B}T\approx J_{4}/k_{\rm B}T\approx 1$, is acceptably nonlinear. Under the same conditions, the variance $\sigma^{2}$ of the equilibrium activation function, panel (b), is generally comparable in scale to the mean of the function.
II.2 Nonequilibrium behavior
@@
Having assessed the equilibrium behavior of the thermodynamic neuron of Fig. 1(a), we turn to its dynamical behavior. Thermodynamic computers operate under Langevin dynamics, both overdamped and underdamped Aifer et al. (2024); Melanson et al. (2023). In this paper we will simulate the behavior of a thermodynamic computer using overdamped Langevin dynamics, in which case our single thermodynamic neuron evolves according to the equation Van Kampen (1992)
$\dot{x}=-\mu\frac{\partial}{\partial x}U_{\bm{J}}(x,I)+\sqrt{2\mu k_{\rm B}T}\,\eta(t),$ (5)
@@
where $U_{\bm{J}}(x,I)$ is given by Eq. (1). Here $\mu$, the mobility parameter, sets the basic time constant of the computer. For the thermodynamic computers of Refs. Aifer et al. (2024); Melanson et al. (2023), $\mu^{-1}\sim 1$ microsecond. For damped oscillators made from mechanical elements Dago et al. (2021) or Josephson junctions Ray and Crutchfield (2023), the time constant would be of order a millisecond or a nanosecond, respectively. The second term on the right-hand side of Eq. (5) represents the thermal fluctuations of the heat bath; $\eta$ is a Gaussian white noise satisfying $\langle\eta(t)\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$.
@@
In Fig. 3(a)–(c) we show the dynamical evolution of the neuron $x$ under the dynamics (5), starting from $x=0$, for various fixed values of the input $I$. Here $\beta=100$. Panel (a) shows the evolution of (5) for the case of a purely quadratic coupling, ${\bm{J}}=(1,0,0)$. Time traces for 11 values of $I$ are shown, evenly spaced from $-5$ to 5. The neuron is initially out of equilibrium, converging to equilibrium in about 2 time units (time is expressed in units of $\mu^{-1}$). As described by Fig. 1(b), the mean value of $x$ in equilibrium is a linear function of $I$. As described by Fig. 2(a), the size of the neuron’s fluctuations in equilibrium is independent of $I$.
@@
Fig. 3(b) shows the case of a purely quartic coupling, ${\bm{J}}=(0,0,1)$. Again, time traces for 11 values of $I$ are shown, evenly spaced from $-5$ to 5; these converge at sufficiently long times to a nonlinear function $x(I)$. As described by Fig. 2(a), the fluctuations of $x$ are largest for $I=0$.
@@
Fig. 3(c) shows the case of a mixed quadratic-quartic coupling, ${\bm{J}}=(1,0,1)$. The long-time mean activation function is nonlinear in $I$, and, as described by Fig. 2(a), the fluctuations near $I=0$ are suppressed relative to the purely quartic case of panel Fig. 3(b).
@@
Figure 4: (a) Bistable neuron potential (blue) at zero input, and (b) the resulting neuron dynamics (5) for $\beta=5$.
@@
From these plots we see that for short times (here less than about 2 time units) the thermodynamic neuron’s finite-time activation function does not equal the equilibrium (long-time) activation function. In Fig. 3(d) we show the value of $x$, derived from Eq. (5), for a range of values of the input $I$, for various fixed values of $t$ (here we set $\beta=\infty$). For times shorter than about 0.2 units, the finite-time activation function of the neuron is linear on the scale of the inputs shown. Thus a network of such neurons would not function as a universal approximator if observed on such timescales. For longer times, however, the finite-time activation function of the neuron is nonlinear on the scale of the inputs shown. As long as we observe the output of a thermodynamic neural network on timescales longer than this threshold, it will behave as a nonlinear model. As observation time increases, the finite-time activation function of the thermodynamic neuron converges to the long-time equilibrium activation function of the neuron, which we have designed to be nonlinear.
@@
For certain values of the intrinsic couplings ${\bm{J}}$, a thermodynamic neuron can display multistability. For example, the potential (1) with ${\bm{J}}=(-|J_{2}|,0,J_{4})$ displays two minima at activations $x_{0}=\pm\sqrt{J_{2}^{2}/(2J_{4})}$, separated by a potential energy barrier of size $J_{2}^{2}/(4J_{4})$. One example of such a potential is shown in Fig. 4(a), compared with the quadratic potential (gray). The resulting dynamics (5) at zero input shows intermittency, as shown in Fig. 4(b). Bistable neurons possess nonlinear activation functions in the vicinity of their stable states, but can also exhibit abrupt changes between states. Such bistability could be used for robust memory storage, or as a mechanism for noise-robust classification in thermodynamic neural networks. In the remainder of this paper we consider only monostable neurons.
@@
Figure 5: Elements of a thermodynamic computer analogous to a neural network. (a) The thermodynamic neurons described in Figs. 13 are connected by bilinear couplings. (b) We consider layered networks of such neurons (adjacent layers are coupled all-to-all), with total potential energy (8).
@@
To summarize the design considerations of this section: 1) the thermodynamic neuron of Fig. 1(a) has a nonlinear equilibrium activation function if it possesses a quartic nonlinearity; 2) it is useful to also include a quadratic nonlinearity in order to suppress equilibrium fluctuations near zero neuron input; and 3) for observation times longer than some threshold, such a neuron also possesses a finite-time activation function that is nonlinear. As a result, networks of such neurons can serve as universal function approximators, both in and out of equilibrium. In the following section we illustrate these properties.
III Programming a thermodynamic computer
III.1 A digital model of a thermodynamic universal approximator
@@
Figure 6: Training a simulation model of a thermodynamic computer to express a nonlinear function at a specified observation time. (a) Loss (14) as a function of evolutionary time $n$ for a layered thermodynamic computer (see Fig. 5) with quadratic neurons (gray) or quadratic-quartic neurons (green). (b) Output (15) at observation time $t_{\rm f}=1$ of the linear computer (gray) and the nonlinear computer (green), as a function of the input $z$, averaged over $M=10^{3}$ samples. The target function (13) is shown as a black line. (c,d). Mean neuron activations measured at observation time $t_{\rm f}$ as a function of the neuron inputs at the same time, for (c) the nonlinear model and (d) the linear model. The color bands denote $\pm$ one standard deviation. Insets: the measured activation at $t_{\rm f}$ compared with the equilibrium neuron activation (2).
@@
$V_{\rm int}({\bm{x}})=\sum_{i=1}^{N}U_{\bm{J}}(x_{i},b_{i})+\sum_{{\rm conn}(ij)}J_{ij}x_{i}x_{j}.$ (6)
The subscript ‘int’ stands for ‘internal’, and ‘conn’ for ‘connections’. The first sum in (6) runs over $N$ single-neuron energy terms (1), with the intrinsic couplings of each neuron set by the vector ${\bm{J}}=(J_{2},J_{3},J_{4})$. Our default choice is $(1,0,1)$. We will also consider the linear-model case $(1,0,0)$, in order to demonstrate the difference in expressive power between a linear model and a nonlinear one. The input $b_{i}$ to each neuron serves as a bias.
The second sum in (6) runs over all distinct pairs of connected neurons, which are determined by the graph structure imposed. We use the bilinear interaction of Refs. Aifer et al. (2024); Melanson et al. (2023). The computers described in those papers use an all-to-all coupling; here, to make contact with existing neural-network designs, we consider the layered structure shown in Fig. 5, with all-to-all connections between layers. This design mimics that of a conventional deep fully-connected neural network. However, unlike in a conventional deep neural network, in which information flows from the input layer to the output layer, the bilinear interaction $J_{ij}x_{i}x_{j}$ ensures that neuron $i$ communicates with neuron $j$, and vice versa, and so information flows forward and backward between the layers of the thermodynamic computer.
@@
$\dot{x}_{i}=-\mu\frac{\partial V({\bm{x}})}{\partial x_{i}}+\sqrt{2\mu k_{\rm B}T}\,\eta_{i}(t),$ (9)
where $V({\bm{x}})$ is given by Eq. (8). The Gaussian white noise terms satisfy $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$.
@@
$\{x_{i}({\bm{I}},t)\},i\in{\rm outputs}.$ (10)
Here ${\bm{I}}$ is the vector of inputs, and $x_{i}({\bm{I}},t)$ denotes the outcome of the dynamics (9) for neuron $i$ at time $t$ upon starting from zero neuron activations, ${\bm{x}}={\bm{0}}$ (all dynamical trajectories start from zero neuron activations).
Because the computer is noisy, we wish to take $M$ samples of each output and average over these samples. We will consider two types of sampling. The first is reset sampling. In this mode of operation we run the computer for time $t_{\rm f}$, observe the outcome, reset the neuron activations to zero, and repeat the procedure $M-1$ times, gathering $M$ samples in total. The advantage of this mode of sampling is that it naturally lends itself to parallelization: the $M$ samples can be computed independently, on distinct copies of the thermodynamic computer if such copies are available. Reset sampling can also be done using a single computer whose neurons are reset periodically. The reset-sampling average is
where the sum runs over $M$ independent realizations $\alpha$ of the dynamics (9). The only requirement on $t_{\rm f}$ is that it is long enough that the effective activation function of the neuron is nonlinear. From the considerations of Section II.2, we set $t_{\rm f}=1$ (in units of $\mu^{-1}$).
@@
$\langle x_{i}({\bm{I}})\rangle_{\rm s}=M^{-1}\sum_{\alpha=1}^{M}x_{i}({\bm{I}},\alpha t_{\rm obs}),$ (12)
@@
where the sum runs over $M$ samples within a single trajectory. Serial sampling is used in equilibrium in Refs. Aifer et al. (2024); Melanson et al. (2023). In equilibrium, a burn-in time is required for the computer to attain equilibrium, and the observation time must be long enough to obtain uncorrelated equilibrium samples. Here, by contrast, we do not require that any portion of the trajectory correspond to equilibrium (it may do, but it is not required to). As a result, we require only that $t_{\rm obs}$ be long enough for the model to be nonlinear (from the considerations of Section II.2, we set $t_{\rm obs}=0.2$, in units of $\mu^{-1}$). No burn-in time is required, and samples do not need to be uncorrelated.
@@
Figure 7: (a–d) Output (15) at time $t_{\rm f}=1$ of the trained nonlinear thermodynamic computer as a function of input $z$, computed using $M$ samples. The target function (13) is shown as a black line (training was done using $M=10^{3}$ samples). (e) Loss (14) as a function of $M$.
@@
Figure 8: (a–d) Output (15) at various observation times $t_{\rm f}$ of the trained nonlinear thermodynamic computer, as a function of input $z$, computed using $M=10^{3}$ samples. The target function (13) is shown as a black line. The computer is trained so that it reproduces the target function when observed at time $t_{\rm f}=1$. (e) Loss (14) as a function of observation time $t_{\rm f}$.
III.2 Programming the computer
@@
In what follows we consider the adjustable parameters of the computer to be the set ${\bm{\theta}}=\{W_{ij}\}\cup\{b_{i}\}\cup\{J_{ij}\}\cup\{f_{i}\}$. Here $\{W_{ij}\}$ is the set of input weights specified by Eq. (7); $\{b_{i}\}$ is the set of biases specified by Eq. (6); $\{J_{ij}\}$ is the set of connections specified by the same equation; and $\{f_{i}\}$ is a set of weights that couple to the output neurons (see Section III.3).
@@
Figure 9: Training a simulation model of a thermodynamic computer to express a nonlinear function in serial-sampling mode. (a) Loss (14) as a function of evolutionary time $n$ for a layered thermodynamic computer (see Fig. 5) with quadratic-quartic $(1,0,1)$ neurons. (b) Output (15) of the thermodynamic computer, evaluated using $M=10^{3}$ time slices, evenly spaced from $t_{\rm obs}=0.2$ to time $t_{\rm f}=Mt_{\rm obs}=200$ (see Eq. (12)). The black line is the target function $y_{0}(z)$, Eq. (13). (c) Color maps of the output $y(z)$ of the trained computer, evaluated using $M=10^{3}$ evenly-spaced time slices up to time $t_{\rm f}$. The bottom panel shows the difference between the computer output and the target function. The steady-state output of the computer remains close to the target function for a range of observation times. (d) Analogous plots for the computer trained in reset-sampling mode (see Figs. 68); here, the computer approximates the target function only at the observation time specified in training ($t_{\rm f}=1$).

### v3

@@ Abstract
$\langle\cdot\rangle_{0}\equiv\frac{\int{\rm d}x\,(\cdot)\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}{\int{\rm d}x\,{\rm e}^{-\beta U_{\bm{J}}(x,I)}}.$ (S3)
When the neuron potential is purely quadratic, i.e. ${\bm{J}}=(J_{2},0,0)$ – with $J_{2}>0$ to ensure thermodynamic stability – the integrals in (S3) can be solved analytically, giving the linear form $m=I/(2J_{2})$. This form is plotted, for $J_{2}=1$, as a gray line in Fig. S1(b); the horizontal dotted black line denotes the value zero. In this case the equilibrium activation function of the neuron is linear, meaning that networks of such neurons in equilibrium cannot serve as universal function approximators  [^4] The quadratic nonlinearity also gives rise to a linear activation function out of equilibrium..
Figure S2: (a) Equilibrium fluctuations (S4) of the thermodynamic neuron of Fig. S1(a), for the case $\beta=100$. The addition of the quadratic coupling to the quartic one (green), suppresses fluctuations relative to the pure quartic case (blue). The mean equilibrium activation functions of those two cases are similar; see Fig. S1(c). The inset shows the largest fluctuations of the purely quartic neuron to be many times that of the purely quadratic neuron. (b,c) The mean (b) and variance (c) of the equilibrium activation function of the $(1,0,1)$ neuron depend on temperature.
Figure S3: Nonequilibrium properties of the thermodynamic neuron of Fig. S1(a). (a–c) Langevin evolution (S5) of the neuron $x$ for $\beta=100$, for 11 evenly-spaced values of $I$. We consider three different sets of couplings $\bm{J}$}: (a) purely quadratic; (b) purely quartic; and (c) mixed quadratic-quartic. (d) The finite-time activation function of the (1,0,1) neuron under the dynamics (S5) (here for $\beta=\infty$) is nonlinear above some threshold observation time (here about 0.2 time units), and for longer times converges to the equilibrium result (curved black dashed line).
The simplest case that is thermodynamically stable and admits a nonlinear equilibrium activation function is the purely quartic case, ${\bm{J}}=(0,0,J_{4})$, with $J_{4}>0$. This case is shown as a blue line in Fig. S1(b), with $J_{4}=1$  [^5] In the purely quartic case the equilibrium activation function can be expressed analytically, in terms of the hypergeometric function.. The equilibrium activation function is nonlinear: its gradient (shown inset) is largest near the origin, and decreases as $|I|$ becomes large. The gradient remains finite for finite $I$: the neuron does not saturate.
Panels (c), (d), and (e) of Fig. S1 show the effect on the thermodynamic neuron’s equilibrium activation function of including quadratic and cubic terms with the quartic coupling. The resulting activation functions display a range of forms that resemble some of those used in classical neural networks, notably the sigmoid and hyperbolic tangent functions LeCun et al. (2015); Schmidhuber (2015); Dubey et al. (2022).
@@
$\sigma^{2}=\langle x^{2}\rangle_{0}-\langle x\rangle_{0}^{2}.$ (S4)
In Fig. S2(a) we plot the value of (S4) as a function of neuron input $I$ for the quadratic case ${\bm{J}}=(1,0,0)$ (gray) and the quartic case ${\bm{J}}=(0,0,1)$ (blue). In the quadratic case the variance is constant for constant temperature, $\sigma^{2}=1/(2\beta)$, reflecting the equipartition theorem. In the quartic case the variance is not constant, and is largest near $I=0$; the inset shows that its maximum value is several times that of the quadratic neuron. The mixed quadratic-quartic case ${\bm{J}}=(1,0,1)$, shown green in Fig. S1(c), has fluctuations at the origin comparable to the quadratic case, while the mean activation function of the case $(1,0,1)$, shown in Fig. S1(c), is similar to that of the pure quartic case. The addition of the quadratic term to the quartic one suppresses fluctuations without changing the essence of the nonlinearity.
@@
The equilibrium activation function of a nonlinear thermodynamic neuron depends on temperature. In Fig. S2(b) and (c) we show the mean and variance of the equilibrium activation function of the neuron ${\bm{J}}=(1,0,1)$ for a range of values of $\beta$. The mean activation is nonlinear provided that temperature is not much larger than the scales of $J_{2}$ and $J_{4}$. The case in which they are comparable, i.e. $|J_{2}|/k_{\rm B}T\approx J_{4}/k_{\rm B}T\approx 1$, is acceptably nonlinear. Under the same conditions, the variance $\sigma^{2}$ of the equilibrium activation function, panel (b), is generally comparable in scale to the mean of the function.
S2.2 Nonequilibrium behavior
@@
Having assessed the equilibrium behavior of the thermodynamic neuron of Fig. S1(a), we turn to its dynamical behavior. Thermodynamic computers operate under Langevin dynamics, both overdamped and underdamped Aifer et al. (2024); Melanson et al. (2025). In this paper we will simulate the behavior of a thermodynamic computer using overdamped Langevin dynamics, in which case our single thermodynamic neuron evolves according to the equation Van Kampen (1992)
$\dot{x}=-\mu\frac{\partial}{\partial x}U_{\bm{J}}(x,I)+\sqrt{2\mu k_{\rm B}T}\,\eta(t),$ (S5)
@@
where $U_{\bm{J}}(x,I)$ is given by Eq. (S1). Here $\mu$, the mobility parameter, sets the basic time constant of the computer. For the thermodynamic computers of Refs. Aifer et al. (2024); Melanson et al. (2025), $\mu^{-1}\sim 1$ microsecond. For damped oscillators made from mechanical elements Dago et al. (2021) or Josephson junctions Ray and Crutchfield (2023), the time constant would be of order a millisecond or a nanosecond, respectively. The second term on the right-hand side of Eq. (S5) represents the thermal fluctuations of the heat bath; $\eta$ is a Gaussian white noise satisfying $\langle\eta(t)\rangle=0$ and $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$.
@@
In Fig. S3(a)–(c) we show the dynamical evolution of the neuron $x$ under the dynamics (S5), starting from $x=0$, for various fixed values of the input $I$. Here $\beta=100$. Panel (a) shows the evolution of (S5) for the case of a purely quadratic coupling, ${\bm{J}}=(1,0,0)$. Time traces for 11 values of $I$ are shown, evenly spaced from 5-5 to 5. The neuron is initially out of equilibrium, converging to equilibrium in about 2 time units (time is expressed in units of $\mu^{-1}$). As described by Fig. S1(b), the mean value of $x$ in equilibrium is a linear function of $I$. As described by Fig. S2(a), the size of the neuron’s fluctuations in equilibrium is independent of $I$.
@@
Fig. S3(b) shows the case of a purely quartic coupling, ${\bm{J}}=(0,0,1)$. Again, time traces for 11 values of $I$ are shown, evenly spaced from 5-5 to 5; these converge at sufficiently long times to a nonlinear function $x(I)$. As described by Fig. S2(a), the fluctuations of $x$ are largest for $I=0$.
@@
Fig. S3(c) shows the case of a mixed quadratic-quartic coupling, ${\bm{J}}=(1,0,1)$. The long-time mean activation function is nonlinear in $I$, and, as described by Fig. S2(a), the fluctuations near $I=0$ are suppressed relative to the purely quartic case of panel Fig. S3(b).
@@
Figure S4: (a) Bistable neuron potential (blue) at zero input, and (b) the resulting neuron dynamics (S5) for $\beta=5$.
@@
From these plots we see that for short times (here less than about 2 time units) the thermodynamic neuron’s finite-time activation function does not equal the equilibrium (long-time) activation function. In Fig. S3(d) we show the value of $x$, derived from Eq. (S5), for a range of values of the input $I$, for various fixed values of $t$ (here we set $\beta=\infty$). For times shorter than about 0.2 units, the finite-time activation function of the neuron is linear on the scale of the inputs shown. Thus a network of such neurons would not function as a universal approximator if observed on such timescales. For longer times, however, the finite-time activation function of the neuron is nonlinear on the scale of the inputs shown. As long as we observe the output of a thermodynamic neural network on timescales longer than this threshold, it will behave as a nonlinear model. As observation time increases, the finite-time activation function of the thermodynamic neuron converges to the long-time equilibrium activation function of the neuron, which we have designed to be nonlinear.
@@
For certain values of the intrinsic couplings $\bm{J}$}, a thermodynamic neuron can display multistability. For example, the potential (S1) with ${\bm{J}}=(-|J_{2}|,0,J_{4})$ displays two minima at activations $x_{0}=\pm\sqrt{J_{2}^{2}/(2J_{4})}$, separated by a potential energy barrier of size $J_{2}^{2}/(4J_{4})$. One example of such a potential is shown in Fig. S4(a), compared with the quadratic potential (gray). The resulting dynamics (S5) at zero input shows intermittency, as shown in Fig. S4(b). Bistable neurons possess nonlinear activation functions in the vicinity of their stable states, but can also exhibit abrupt changes between states. Such bistability could be used for robust memory storage, or as a mechanism for noise-robust classification in thermodynamic neural networks. In the remainder of this paper we consider only monostable neurons.
@@
Figure S5: Elements of a thermodynamic computer analogous to a neural network. (a) The thermodynamic neurons described in Figs. S1S3 are connected by bilinear couplings. (b) We consider layered networks of such neurons (adjacent layers are coupled all-to-all), with total potential energy (S8).
@@
To summarize the design considerations of this section: 1) the thermodynamic neuron of Fig. S1(a) has a nonlinear equilibrium activation function if it possesses a quartic nonlinearity; 2) it is useful to also include a quadratic nonlinearity in order to suppress equilibrium fluctuations near zero neuron input; and 3) for observation times longer than some threshold, such a neuron also possesses a finite-time activation function that is nonlinear. As a result, networks of such neurons can serve as universal function approximators, both in and out of equilibrium. In the following section we illustrate these properties.
S3 Programming a thermodynamic computer
S3.1 A digital model of a thermodynamic universal approximator
@@
Figure S6: Training a simulation model of a thermodynamic computer to express a nonlinear function at a specified observation time. (a) Loss (S14) as a function of evolutionary time $n$ for a layered thermodynamic computer (see Fig. S5) with quadratic neurons (gray) or quadratic-quartic neurons (green). (b) Output (S15) at observation time $t_{\rm f}=1$ of the linear computer (gray) and the nonlinear computer (green), as a function of the input $z$, averaged over $M=10^{3}$ samples. The target function (S13) is shown as a black line. (c,d). Mean neuron activations measured at observation time $t_{\rm f}$ as a function of the neuron inputs at the same time, for (c) the nonlinear model and (d) the linear model. The color bands denote $\pm$ one standard deviation. Insets: the measured activation at $t_{\rm f}$ compared with the equilibrium neuron activation (S2).
@@
$V_{\rm int}({\bm{x}})=\sum_{i=1}^{N}U_{\bm{J}}(x_{i},b_{i})+\sum_{{\rm conn}(ij)}J_{ij}x_{i}x_{j}.$ (S6)
The subscript ‘int’ stands for ‘internal’, and ‘conn’ for ‘connections’. The first sum in (S6) runs over $N$ single-neuron energy terms (S1), with the intrinsic couplings of each neuron set by the vector ${\bm{J}}=(J_{2},J_{3},J_{4})$. Our default choice is $(1,0,1)$. We will also consider the linear-model case $(1,0,0)$, in order to demonstrate the difference in expressive power between a linear model and a nonlinear one. The input $b_{i}$ to each neuron serves as a bias.
The second sum in (S6) runs over all distinct pairs of connected neurons, which are determined by the graph structure imposed. We use the bilinear interaction of Refs. Aifer et al. (2024); Melanson et al. (2025). The computers described in those papers use an all-to-all coupling; here, to make contact with existing neural-network designs, we consider the layered structure shown in Fig. S5, with all-to-all connections between layers. This design mimics that of a conventional deep fully-connected neural network. However, unlike in a conventional deep neural network, in which information flows from the input layer to the output layer, the bilinear interaction $J_{ij}x_{i}x_{j}$ ensures that neuron $i$ communicates with neuron $j$, and vice versa, and so information flows forward and backward between the layers of the thermodynamic computer.
@@
$\dot{x}_{i}=-\mu\frac{\partial V({\bm{x}})}{\partial x_{i}}+\sqrt{2\mu k_{\rm B}T}\,\eta_{i}(t),$ (S9)
where $V({\bm{x}})$ is given by Eq. (S8). The Gaussian white noise terms satisfy $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$.
@@
$\{x_{i}({\bm{I}},t)\},i\in{\rm outputs}.$ (S10)
Here $\bm{I}$} is the vector of inputs, and $x_{i}({\bm{I}},t)$ denotes the outcome of the dynamics (S9) for neuron $i$ at time $t$ upon starting from zero neuron activations, ${\bm{x}}={\bm{0}}$ (all dynamical trajectories start from zero neuron activations).
Because the computer is noisy, we wish to take $M$ samples of each output and average over these samples. We will consider two types of sampling. The first is reset sampling. In this mode of operation we run the computer for time $t_{\rm f}$, observe the outcome, reset the neuron activations to zero, and repeat the procedure M1M-1 times, gathering $M$ samples in total. The advantage of this mode of sampling is that it naturally lends itself to parallelization: the $M$ samples can be computed independently, on distinct copies of the thermodynamic computer if such copies are available. Reset sampling can also be done using a single computer whose neurons are reset periodically. The reset-sampling average is
$\langle x_{i}({\bm{I}})\rangle_{\rm r}=M^{-1}\sum_{\alpha=1}^{M}x^{(\alpha)}_{i}({\bm{I}},t_{\rm f}),$ (S11)
where the sum runs over $M$ independent realizations $\alpha$ of the dynamics (S9). The only requirement on $t_{\rm f}$ is that it is long enough that the effective activation function of the neuron is nonlinear. From the considerations of Section S2.2, we set $t_{\rm f}=1$ (in units of $\mu^{-1}$).
@@
$\langle x_{i}({\bm{I}})\rangle_{\rm s}=M^{-1}\sum_{\alpha=1}^{M}x_{i}({\bm{I}},\alpha t_{\rm obs}),$ (S12)
@@
where the sum runs over $M$ samples within a single trajectory. Serial sampling is used in equilibrium in Refs. Aifer et al. (2024); Melanson et al. (2025). In equilibrium, a burn-in time is required for the computer to attain equilibrium, and the observation time must be long enough to obtain uncorrelated equilibrium samples. Here, by contrast, we do not require that any portion of the trajectory correspond to equilibrium (it may do, but it is not required to). As a result, we require only that $t_{\rm obs}$ be long enough for the model to be nonlinear (from the considerations of Section S2.2, we set $t_{\rm obs}=0.2$, in units of $\mu^{-1}$). No burn-in time is required, and samples do not need to be uncorrelated.
@@
Figure S7: (a–d) Output (S15) at time $t_{\rm f}=1$ of the trained nonlinear thermodynamic computer as a function of input $z$, computed using $M$ samples. The target function (S13) is shown as a black line (training was done using $M=10^{3}$ samples). (e) Loss (S14) as a function of $M$.
@@
Figure S8: (a–d) Output (S15) at various observation times $t_{\rm f}$ of the trained nonlinear thermodynamic computer, as a function of input $z$, computed using $M=10^{3}$ samples. The target function (S13) is shown as a black line. The computer is trained so that it reproduces the target function when observed at time $t_{\rm f}=1$. (e) Loss (S14) as a function of observation time $t_{\rm f}$.
S3.2 Programming the computer
@@
In what follows we consider the adjustable parameters of the computer to be the set ${\bm{\theta}}=\{W_{ij}\}\cup\{b_{i}\}\cup\{J_{ij}\}\cup\{f_{i}\}$. Here $\{W_{ij}\}$ is the set of input weights specified by Eq. (S7); $\{b_{i}\}$ is the set of biases specified by Eq. (S6); $\{J_{ij}\}$ is the set of connections specified by the same equation; and $\{f_{i}\}$ is a set of weights that couple to the output neurons (see Section S3.3).
@@
Figure S9: Training a simulation model of a thermodynamic computer to express a nonlinear function in serial-sampling mode. (a) Loss (S14) as a function of evolutionary time $n$ for a layered thermodynamic computer (see Fig. S5) with quadratic-quartic $(1,0,1)$ neurons. (b) Output (S15) of the thermodynamic computer, evaluated using $M=10^{3}$ time slices, evenly spaced from $t_{\rm obs}=0.2$ to time $t_{\rm f}=Mt_{\rm obs}=200$ (see Eq. (S12)). The black line is the target function $y_{0}(z)$, Eq. (S13). (c) Color maps of the output $y(z)$ of the trained computer, evaluated using $M=10^{3}$ evenly-spaced time slices up to time $t_{\rm f}$. The bottom panel shows the difference between the computer output and the target function. The steady-state output of the computer remains close to the target function for a range of observation times. (d) Analogous plots for the computer trained in reset-sampling mode (see Figs. S6S8); here, the computer approximates the target function only at the observation time specified in training ($t_{\rm f}=1$).